Circular No. 5871

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993aa IN ANONYMOUS GALAXY C. Pollas, Observatoire de la Côte d’Azur, reports his discovery of a

supernova of mag V = 17.5-18 located at α = 23^{h}00^{m}46^{s}.40, δ = −6^{o}37′06′′.5 (equinox 1950.0). SN 1993aa, found on 0.9-m Schmidt telescope films taken Sept. 19.13 and 20.02 UT, is superimposed on the galaxy nucleus, the galaxy itself being of m_{pg} ∼ 18 and rather blue on the Palomar Sky Survey. No such object appears at the position of the supernova on either the Palomar Sky Survey or an R film taken on 1991 Sept. 20. A nearby star of mag ∼ 15 is located at α = 23^{h}00^{m}47^{s}.23, δ = −6^{o}39′03′′.5.

A. V. Filippenko and T. Matheson, University of California at Berkeley, report that uncalibrated CCD spectra (range 310-1000 nm) obtained on Sept. 25 UT with the Shane 3-m reflector at Lick Observatory show that the supernova is of type Ia, probably ∼ 1 month past maximum brightness.

SUPERNOVA 1993ab IN NGC 1164 J. Mueller reports the discovery of a supernova (blue mag ∼ 18) lo-

cated 17′′.5 east and 12′′ south of the nucleus of NGC 1164 (α = 2^{h}58^{m}.7, δ = +42^{o}24′, equinox 1950.0). The supernova was found on a plate taken on Sept. 24 UT by J. D. Mendenhall and Mueller with the 1.2-m Oschin Schmidt Telescope in the course of the second Palomar Sky Survey. A spectrogram obtained by S. G. Djorgovski, J. D. Smith, and D. Thompson on Sept. 25 UT with the 5-m Hale Telescope (+ 4-shooter + spectrograph) shows broad emission lines and a blue continuum typical of a type-II supernova. There is a hint that the supernova might be increasing in brightness.

NOVA LUPI 1993 T. Kato, Kyoto University, reports that a pre-discovery image of N

Lup 1993 has been found on Fujichrome P1600 photographs taken by I. Okada, observing from the Yulara Resort, N.T., Australia; Okada reports the following photographic magnitudes: Sept. 15.051 UT, 7.7; 16.465, 7.9; 16.488, 8.0.

Visual magnitude estimates: Sept. 23.32 UT, 8.5 (A. F. Jones, Nelson, New Zealand); 23.72, 8.8 (D. Overbeek, Edenvale, S. Africa).

1993 September 25 Daniel W. E. Green